A quality control analyst mixes 30.0ml of 0.255M BaCl2 with excess AgNo3 causing AgCl to precipitate and subsequently isolated 2.01g of AgCl. Show the balanced equation for the reaction and calculate the percent yield of the reaction


1
Expert's answer
2021-08-24T02:19:48-0400

AgNO 3 (aq) + BaCl 2 (aq) → 2 AgCl (s) + Ba(NO 3) 2 (aq)

Molar mass of BaCl2 = 208.23 g/mol

0.255× 208.23 = 53.09865

53.09865/30 = 1.77

= 1.77/2.01× 100

=88.06%